Determine the pressure exerted by a gas at –50.0°C if that gas exerted a pressure of 1.22 bar when the temperature was 50.0°C.
Answer #1

T = 50€ = 50+273 T= -500 =_50+273 - 2230 a 323k 1.22 = P2 323 P = 0.842 bar